Find two numbers, one of which is double the other, if the square of their sum exceeds the sum of their squares by 1oo.
Answer by Alan3354(69443)

Find two numbers, one of which is double the other, if the square of their sum exceeds the sum of their squares by 1oo.
----------------------
x and 2x
(x+2x)^2 = x^2 + (2x)^2 + 100
9x^2 = 5x^2 + 100
4x^2 = 100
x^2 = 25
x = 5, -5
-----------
The numbers are 5 and 10, or
-5 and -10